# Certification of high stellar ranks of quantum states of light with a pair of click detectors

### v3

@@ IV Stellar rank thresholds
In order to determine the stellar rank thresholds (7) we first write down the POVM element corresponding to click of D1 and no-click of D2,
$\hat{W}=\hat{\Pi}_{1}=\sum_{n=0}^{\infty}R_{1}(n)|n\rangle\langle n|.$ (15)
This explicitly reads
@@
To proceed, we recall the expression for density matrix of a thermal state with mean number of photons $\bar{n}$,
@@
According to the Bloch-Messiah decomposition we can express any single-mode unitary Gaussian operation as a sequence of phase shifts, single-mode suqeezing and coherent displacement [86],
@@
Analytical formulas for density matrix elements $\rho_{G,mn}=\langle m|\hat{\rho}_{G}|n\rangle$ of the displaced squeezed thermal state have been derived in Refs. [87, 88]. For completeness, we recall that these density matrix elements can be calculated using the well-known fact that the Husimi $Q$-function
@@
Here $\beta$ and $\beta^{\ast}$ are formally treated as two independent variables. The displaced squeezed thermal state $\hat{\rho}_{G}$ possesses Gaussian Husimi $Q$-function
@@
and the complex amplitudes are decomposed into their real and imaginary parts, $\alpha=\alpha_{R}+i\alpha_{I}$, and $\beta=\beta_{R}+i\beta_{I}$. On inserting the expression (25) into formula (24) one obtains [88]
@@
×$\times\sum_{j=0}^{\min(m,n)}\frac{A^{m+n-2j}B^{j}}{j!(m-j)!(n-j)!}H_{m-j}(\omega)H_{n-j}(\omega^{\ast}).$
Here $H_{n}(\cdot)$ denotes Hermite polynomial of degree $n$, and the parameters $A$, $B$, and $\omega$ read
@@
The optimization of the stellar rank thresholds $W_{m}$ over three real parameters $\alpha_{R}$, $\alpha_{I}$ and $r$ was performed numerically by an exhaustive search. The numerics reveals that it is optimal to displace the state along one of the principal squeezing axes, which is consistent with previous observations [46, 48, 30]. The final optimization was therefore performed over real non-negative $\alpha$ and real $r$ (both positive and negative). Numerical results for a balanced scheme with $T=\frac{1}{2}$ are summarized in Table I. In addition to the stellar rank thresholds we also specify in the table the maximum possible value of the witness, i.e. maximum of $R_{1}(n)$ over all $n$ for given $\eta$. We can see that the maximum stellar rank that can be certified by measurement of the single-click probability $R_{1}$ increases with decreasing detection efficiency $\eta$, which is consistent with the shift of the maximum of $R_{1}(n)$ to higher $n$, c.f. Fig. 2. However, for low $\eta$ and high stellar ranks $m$, the interval $W_{\mathrm{max}}-W_{m}$ becomes very narrow. This indicates that the witness can certify only a narrow subset of states with stellar rank $m$ and the measurement precision would have to be very high to achieve reliable certification. In Table II we present the stellar rank thresholds for several different values of beam splitter transmittance $T$. In accordance with the behavior observed in Fig. 3, then maximum certifiable stellar rank increases with increasing $T$.
@@
Table 1: Stellar rank thresholds $W_{m}$ for the witness $\hat{W}=\hat{\Pi}_{1}$ specified in Eq. (15). The thresholds $W_{m}$ are presented for $T=0.5$ and four different values of the total detection efficiency $\eta$. The witnesses $W_{m}$ are rounded up to four decimal places. The last row specifies the maximum physically achieveable value of the witness.
@@ V Witnesses based on $R_{1}$ and $R_{2}$
Figure 4: Witnesses of stellar rank based on the pair of probabilities $R_{1}$ and $R_{2}$ are plotted for balanced setup with $T=0.5$, and total detection efficiency $\eta=1$ (a) and $\eta=0.5$ (b). The color of each area indicates the minimum stellar rank that is certified when the point $[R_{2},R_{1}]$ lies in that area. Points in the light gray area are compatible with Gaussian states and their mixtures. For $\eta=1$, only stellar rank $1$ can be certified (the light blue area), while higher stellar ranks become certifiable for $\eta=0.5$. The black lines indicate the boundary of all physically achievable probability pairs $[R_{2},R_{1}]$. See text for additional details.
In this section we consider more general class of stellar rank witnesses based on the pair of probabilities of $R_{1}$ and $R_{2}$ where $R_{2}$ denotes the probability that both detectors D1 and D2 click simultaneously [85, 72, 30, 83]. We shall first consider balanced detection scheme with $T=\frac{1}{2}$. The probability $R_{2}(n)$ for input Fock state $n$ is then given by
@@
Here we made use of the identity $R_{2}=1-2R_{1}-R_{0}$ valid for balanced two-detector scheme, and $R_{0}(n)=(1-\eta)^{n}$ is the probability that none of the detectors click. The POVM element $\hat{\Pi}_{2}$ associated with this detector outcome reads
$\hat{\Pi}_{2}=\sum_{n=0}^{\infty}R_{2}(n)|n\rangle\langle n|.$ (30)
Note that $\hat{\Pi}_{2}$ can be expressed as a linear combination of identity operator and density matrices of two thermal states.
Following the long-established approach of Refs. [46, 85, 30, 72], we can consider witnesses formed by linear combinations of $\hat{\Pi}_{1}$ and $\hat{\Pi}_{2}$,
$\hat{W}_{\lambda}=\hat{\Pi}_{1}-\lambda\hat{\Pi}_{2}.$ (31)
@@
The set of probability pairs $[R_{2},R_{1}]$ that are achievable by quantum states with stellar rank at most $m-1$ forms a convex set $\mathcal{S}_{m-1}$ in a two-dimensional space, see Fig. 4. Determination of the stellar rank threshold $W_{\lambda,m}$ for the witness $\hat{W}_{\lambda}$ identifies a line that is tangent to the boundary of this convex set at some point. By varying the parameter $\lambda$ we can probe and find the whole boundary of this convex set $\mathcal{S}_{m-1}$.
For reference, we first plot in Fig. 4(a) the results for $\eta=1$ [30]. In this case, the setup allows to certify only the stellar rank $1$. Note that analytical description of the quantum non-Gaussianity boundary depicted in Fig. 4(a) was derived in Ref. [48]. Besides the boundary between the various sets $\mathcal{S}_{m}$ it is also important to identify the ultimate boundary of all physically achievable probability pairs $[R_{2},R_{1}]$. For $\eta=1$, this boundary is formed by a triangle [48], as illustrated in Fig. 4(a). The vertices of this triangle correspond to the vacuum state $[0,0]$, single-photon Fock state $[0,\frac{1}{2}]$, and a Gaussian quantum state with asymptotically infinite energy so that both detectors always click, $[1,0]$. Note that the probability pairs $[R_{2}(n),R_{1}(n)]$ for input Fock states with $n>1$ all lie on a straight line that connects the points $[0,\frac{1}{2}]$ and $[1,0]$. To show this, note that for $\eta=1$ we get $R_{2}(n)=1-2/2^{n}+\delta_{n0}$ and $R_{1}(n)=1/2^{n}-\delta_{n0}$. For $n>0$ we thus obtain a simple linear relationship
@@
As illustrated in Fig. 4(b), this picture changes for $\eta<1$ and higher stellar ranks become certifiable for certain states. Note that the witnesses based on pair of probabilities $R_{1}$ and $R_{2}$ are strictly stronger than witness based just on $R_{1}$, because the knowledge of probability of simultaneous clicks of both detectors $R_{2}$ provides additional information that imposes stricter limit on the maximum achievable probability $R_{1}$ of click of the first detector only. Note also that the quantum non-Gaussianity criterion for this type of lossy detection with a pair of click detectors was presented in Ref. [83], c.f. the boundary of the gray area in Fig. 4(b) .
For $\eta<1$ the boundary of the set of all physically achievable probability pairs $[R_{2},R_{1}]$ is not a simple triangle but is instead formed by an infinite number of straight lines that connect the points $[R_{2}(n),R_{1}(n)]$ and $R_{2}(n+1),R_{1}(n+1)]$. The probability pairs $[R_{2}(n),R_{1}(n)]$ corresponding to various Fock states $|n\rangle$ do not lie on a single straight line anymore but on a concave curve (see Appendix for more details). The boundary of the convex hull of this set is thus formed by the above specified sequence of straight lines. This structure of the physical boundary suggests that stellar rank $m$ of arbitrary high Fock state $|m\rangle$ is in principle certifiable by this scheme. However, for large $m$ the boundaries of the various sets $\mathcal{S}_{m}$ practically coincide. Successful application of the witness to high Fock states may therefore require extremely high precision of the measurement of click probabilities $R_{1}$ and $R_{2}$.
Figure 5: The same as Fig. 4, but $T=0.75$. Witnesses of stellar rank based on the pair of probabilities $R_{1}$ and $R_{2}$ are plotted for total detection efficiency $\eta=1$ (a) and $\eta=0.5$ (b). The color of each area indicates the minimum stellar rank that is certified when the point $[R_{2},R_{1}]$ lies in that area. The black lines indicate the boundary of all physically achievable probability pairs $[R_{2},R_{1}]$.
@@
and the corresponding POVM element $\hat{\Pi}_{2}$ can be written as a linear combination of identity operator and density matrices of three thermal states,
$\hat{\Pi}_{2}=\hat{I}-(1+\bar{n}_{1})\rho_{\mathrm{th}}(\bar{n}_{1})-(1+\bar{n}_{3})\rho_{\mathrm{th}}(\bar{n}_{3})+(1+\bar{n}_{2})\rho_{\mathrm{th}}(\bar{n}_{2}),$ (35)
where $\bar{n}_{1}=1/[\eta(1-T)]-1$, $\bar{n}_{2}=1/\eta-1$, and $\bar{n}_{3}=1/(\eta T)-1$, see also Eq. (19).
@@
In Figure 5 we plot the numerical results for $T=0.75$. We can see that for unbalanced detection scheme the witness based on the pair of probabilities of click of the first detector only $R_{1}$ and probability of simultaneous click click of both detectors $R_{2}$ can detect stellar ranks higher than $1$ even for $\eta=1$. However, the area corresponding to detection of higher stellar ranks is very small. When the detection efficiency is decreased to $\eta=0.5$ then also the area corresponding to the certification of stellar rank $3$ becomes visible, see Fig. 5(b).
VI Conclusions

### v4

@@ IV Stellar rank thresholds
In order to determine the stellar rank thresholds (4) we first write the POVM element corresponding to click of D1 and no-click of D2:
$\hat{W}=\hat{R}_{1}=\sum_{n=0}^{\infty}R_{1}(n)|n\rangle\langle n|.$ (22)
This explicitly reads as
@@
To proceed, we recall the expression for density matrix of a thermal state with mean number of photons $\bar{n}$:
@@
According to the Bloch-Messiah decomposition we can express any single-mode unitary Gaussian operation as a sequence of phase shifts, single-mode suqeezing, and coherent displacement [88],
@@
Analytical formulas for density matrix elements $\rho_{G,mn}=\langle m|\hat{\rho}_{G}|n\rangle$ of the displaced squeezed thermal state have been derived in Refs. [89, 90]. For completeness, we recall that these density matrix elements can be calculated using the well-known fact that the Husimi $Q$ function
@@
Here $\beta$ and $\beta^{\ast}$ are formally treated as two independent variables. The displaced squeezed thermal state $\hat{\rho}_{G}$ possesses Gaussian Husimi $Q$ function
@@
and the complex amplitudes are decomposed into their real and imaginary parts, $\alpha=\alpha_{R}+i\alpha_{I}$ and $\beta=\beta_{R}+i\beta_{I}$. On inserting the expression (32) into formula (31) one obtains [90]
@@
$\times\sum_{j=0}^{\min(m,n)}\frac{A^{m+n-2j}B^{j}}{j!(m-j)!(n-j)!}H_{m-j}(\omega)H_{n-j}(\omega^{\ast}).$
Here $H_{n}(\cdot)$ denotes Hermite polynomial of degree $n$, and the parameters $A$, $B$, and $\omega$ read as
@@
The optimization of the stellar rank thresholds $W_{m}$ over three real parameters $\alpha_{R}$, $\alpha_{I}$, and $r$ was performed numerically by an exhaustive search. The numerics reveals that it is optimal to displace the state along one of the principal squeezing axes, which is consistent with previous observations [46, 48, 30]. The final optimization was therefore performed over real non-negative $\alpha$ and real $r$ (both positive and negative). Numerical results for a balanced scheme with $T=\frac{1}{2}$ are summarized in Table I. In addition to the stellar rank thresholds we also specify in the table the maximum possible value of the witness, i.e., maximum of $R_{1}(n)$ over all $n$ for given $\eta$. We can see that the maximum stellar rank that can be certified by measurement of the single-click probability $R_{1}$ increases with decreasing detection efficiency $\eta$, which is consistent with the shift of the maximum of $R_{1}(n)$ to higher $n$ [cf. Fig. 2 and Eq. (18)]. However, the interval $W_{\mathrm{max}}-W_{m}$ becomes very narrow as $m$ increases. This indicates that the witness can certify only a narrow subset of states with stellar rank $m$ and the measurement precision would have to be very high to achieve reliable certification. In Table II we present the stellar rank thresholds for several different values of beam-splitter transmittance $T$. In accordance with the behavior observed in Fig. 3, then maximum certifiable stellar rank increases with increasing $T$.
@@
$n_{\mathrm{max}}$ 1 2 3 6
Table 1: Stellar rank thresholds $W_{m}$ for the witness $\hat{W}=\hat{R}_{1}$ specified in Eq. (22). The thresholds $W_{m}$ are presented for $T=0.5$ and four different values of the total detection efficiency $\eta$. The witnesses $W_{m}$ are rounded up to four decimal places. The last two rows specifiy the maximum physically achievable value of the witness and the Fock number $n_{\mathrm{max}}$ for which $R_{1}(n_{\mathrm{max}})=W_{\mathrm{max}}$ is reached.
@@ V Witnesses based on $R_{1}$ and $R_{2}$
Figure 4: Witnesses of stellar rank based on the pair of probabilities $R_{1}$ and $R_{2}$ are plotted for balanced setup with $T=0.5$, and total detection efficiency $\eta=1$ (a) and $\eta=0.5$ (b). The color of each area indicates the minimum stellar rank that is certified when the point $R_{2},R_{1}]$ lies in that area. Points in the light gray area are compatible with Gaussian states and their mixtures. For $\eta=1$, only stellar rank $1$ can be certified (the light blue area), while stellar ranks higher than one become certifiable for $\eta=0.5$. The black lines indicate the boundary of all physically achievable probability pairs $R_{2},R_{1}]$. See text for additional details.
In this section we consider more general class of stellar rank witnesses based on the pair of probabilities of $R_{1}$ and $R_{2}$, where $R_{2}$ denotes the probability that both detectors D1 and D2 click simultaneously [87, 72, 30, 83]. We shall first consider balanced detection scheme with $T=\frac{1}{2}$. The probability $R_{2}(n)$ for input Fock state $n$ is then given by
@@
Here we made use of the identity $R_{2}=1-2R_{1}-R_{0}$ valid for balanced two-detector scheme, and $R_{0}(n)=(1-\eta)^{n}$ is the probability that none of the detectors click for input Fock state $|n\rangle$. The POVM element $\hat{R}_{2}$ associated with this detector outcome reads as
$\hat{R}_{2}=\sum_{n=0}^{\infty}R_{2}(n)|n\rangle\langle n|.$ (37)
Note that $\hat{R}_{2}$ can be expressed as a linear combination of identity operator and density matrices of two thermal states.
Following the long-established approach of Refs. [46, 87, 30, 72], we can consider witnesses formed by linear combinations of $\hat{R}_{1}$ and $\hat{R}_{2}$,
$\hat{W}_{\lambda}=\hat{R}_{1}-\lambda\hat{R}_{2}.$ (38)
@@
The set of probability pairs $R_{2},R_{1}]$ that are achievable by quantum states with stellar rank at most m1m-1 forms a convex set $\mathcal{S}_{m-1}$ in a two-dimensional space, see Fig. 4. Determination of the stellar rank threshold $W_{\lambda,m}$ for the witness $\hat{W}_{\lambda}$ identifies a line that is tangent to the boundary of this convex set at some point. By varying the parameter $\lambda$ we can probe and find the whole boundary of this convex set $\mathcal{S}_{m-1}$.
For reference, we first plot in Fig. 4(a) the results for $\eta=1$ [30]. In this case, the setup allows to certify only the stellar rank $1$. Note that analytical description of the quantum non-Gaussianity boundary depicted in Fig. 4(a) was derived in Ref. [48]. Aside from the boundary between the various sets $\mathcal{S}_{m}$ it is also important to identify the ultimate boundary of all physically achievable probability pairs $R_{2},R_{1}]$. For $\eta=1$, this boundary is formed by a triangle [48], as illustrated in Fig. 4(a). The vertices of this triangle correspond to the vacuum state $[0,0]$, single-photon Fock state $0,\frac{1}{2}]$, and a Gaussian quantum state with asymptotically infinite energy so that both detectors always click, $[1,0]$. Note that the probability pairs $R_{2}(n),R_{1}(n)]$ for input Fock states with $n>1$ all lie on a straight line that connects the points $0,\frac{1}{2}]$ and $[1,0]$. To show this, note that for $\eta=1$ we get $R_{2}(n)=1-2/2^{n}+\delta_{n0}$ and $R_{1}(n)=1/2^{n}-\delta_{n0}$. For $n>0$ we thus obtain a simple linear relationship
@@
As illustrated in Fig. 4(b), this picture changes for $\eta<1$ and stellar ranks higher than $1$ become certifiable for certain states. Note that the witnesses based on pair of probabilities $R_{1}$ and $R_{2}$ are strictly stronger than witness based just on $R_{1}$, because the knowledge of probability of simultaneous clicks of both detectors $R_{2}$ provides additional information that imposes stricter limit on the maximum achievable probability $R_{1}$ of click of the first detector only. Note also that the quantum non-Gaussianity criterion for this type of lossy detection with a pair of click detectors was presented in Ref. [83] [cf. the boundary of the gray area in Fig. 4(b)].
For $\eta<1$ the boundary of the set of all physically achievable probability pairs $R_{2},R_{1}]$ is not a simple triangle but is instead formed by an infinite number of straight lines that connect the points $R_{2}(n),R_{1}(n)]$ and $R_{2}(n+1),R_{1}(n+1)]$. The probability pairs $R_{2}(n),R_{1}(n)]$ corresponding to various Fock states $|n\rangle$ do not lie on a single straight line anymore but on a concave curve (see Appendix for more details). The boundary of the convex hull of this set is thus formed by the above specified sequence of straight lines. This structure of the physical boundary suggests that stellar rank $m$ of arbitary high Fock state $|m\rangle$ is in principle certifiable by this scheme. However, for large $m$ the boundaries of the various sets $\mathcal{S}_{m}$ practically coincide. Successful application of the witness to high Fock states may therefore require extremely high precision of the measurement of click probabilities $R_{1}$ and $R_{2}$.
Figure 5: The same as Fig. 4, but $T=0.75$. Witnesses of stellar rank based on the pair of probabilities $R_{1}$ and $R_{2}$ are plotted for total detection efficiency $\eta=1$ (a) and $\eta=0.5$ (b). The color of each area indicates the minimum stellar rank that is certified when the point $R_{2},R_{1}]$ lies in that area. The black lines indicate the boundary of all physically achievable probability pairs $R_{2},R_{1}]$.
@@
and the corresponding POVM element $\hat{R}_{2}$ can be written as a linear combination of identity operator and density matrices of three thermal states,
$\hat{R}_{2}=\hat{I}-(1+\bar{n}_{1})\rho_{\mathrm{th}}(\bar{n}_{1})-(1+\bar{n}_{3})\rho_{\mathrm{th}}(\bar{n}_{3})+(1+\bar{n}_{2})\rho_{\mathrm{th}}(\bar{n}_{2}),$ (42)
where $\bar{n}_{1}=1/[\eta(1-T)]-1$, $\bar{n}_{2}=1/\eta-1$, and $\bar{n}_{3}=1/(\eta T)-1$ [see also Eq. (26)].
@@
In Figure 5 we plot the numerical results for $T=0.75$. We can see that for unbalanced detection scheme the witness based on the pair of probabilities of click of the first detector only $R_{1}$ and probability of simultaneous click click of both detectors $R_{2}$ can detect stellar ranks higher than $1$ even for $\eta=1$. However, the area corresponding to detection of higher stellar ranks is very small. When the detection efficiency is decreased to $\eta=0.5$ then also the area corresponding to the certification of stellar rank $3$ becomes visible [see Fig. 5(b)].
VI Experimental considerations
Several experimental factors can influence the ability to certify stellar ranks higher than one with the proposed approach. First of all, the total number of measurements $N$ must be high enough to suppress the statistical errors. To simplify the analysis we shall consider here the stellar rank witnesses based solely on the probability $R_{1}=\sum_{n}d_{n}R_{1}(n)$, where $d_{n}=\langle n|\hat{\rho}|n\rangle$. The required number of measurements to certify stellar rank $m$ can be estimated by requiring that the gap $R_{1}-W_{m}$ between $R_{1}$ and the stellar rank threshold $W_{m}$ is larger than three standard deviations of the estimate of $R_{1}$:
$R_{1}-W_{m}\gtrsim 3\sqrt{\frac{R_{1}(1-R_{1})}{N}}.$ (43)
Formula (43) yields
$N\gtrsim\frac{9R_{1}(1-R_{1})}{(R_{1}-W_{m})^{2}}.$ (44)
Assuming that the actual value of $R_{1}$ is $W_{\mathrm{max}}$ (hence the characterized state is a pure Fock state), and plugging in the data from Tables I and II we find that reliable certification of stellar rank 2 is possible with $N\approx 10^{4}$ samples, while certification of stellar rank $3$ may be possible with slightly less than $10^{5}$ samples. The number of required samples $N$ will further increase for imperfect Fock states and is mainly sensitive to the width of the gap $R_{1}-W_{m}$.
Dark counts represent another factor that can influence the measurement. We consider here a simple model of dark counts that appear with probability $P_{D}$ that does not depend on the state which is measured by the detector. Presence of dark counts modifies the expression for $R_{1}(n)$ as
$R_{1}(n)=(1-P_{D})[1-\eta(1-T)]^{n}-(1-P_{D})^{2}(1-\eta)^{n},$ (45)
and similar generalization can be performed also for $R_{2}(n)$. The stellar rank witnesses and thresholds can be determined for this modified and extended detection model in the same way as for the original model without dark counts. In practice, the dark counts can be of the order of 100 per second or even less. Considering measurement with detection window whose width is several nanoseconds, the probability of dark counts will be $P_{D}\lesssim 10^{-6}$, hence typically negligible.
Finally, the precision of calibration of the parameters $\eta$ and $T$ can also play important role. It may be particularly simple to achieve the balanced operation $T=\frac{1}{2}$, as in this setting the response of the two detectors in Fig. 1 shall be the same (up to statistical fluctuations). Precise absolute calibration of losses may be more challenging. Let us suppose that $\eta$ can be calibrated with mean value $\bar{\eta}$ and residual uncertainty $\Delta\eta$. Then one should consider the stellar rank threshold $W_{m}(\eta)$ as a function of $\eta$ and choose the maximum in the interval of possible values of $\eta$,
$W_{m}=\max_{\eta\in\mathcal{I}}W_{m}(\eta),\qquad\mathcal{I}=[\bar{\eta}-\Delta\eta,\bar{\eta}+\Delta\eta].$ (46)
As can be seen from data in Tables I and II, the stellar rank thresholds typically increase with increasing $\eta$. A conservative choice will therefore usually be to consider the maximum possible value of $\eta$, given by $\bar{\eta}+\Delta\eta$. This choice will conservatively assign part of the actual losses to the incoming state. Increased stellar rank thresholds will lead to increased numbers of required samples $N$, as indicated by Eq. (44). Moreover, if the uncertainty of $\eta$ will be large enough, then the certification of stellar ranks higher than one may become unfeasible. Numerical calculations for $T=0.5$ and $\eta=0.6$ show that the threshold $W_{2}=0.3119$ increases to $W_{2}=0.3144$ for $\Delta\eta=0.01$ and becomes $W_{2}=0.3265$ for $\Delta\eta=0.05$, which is already very close to the maximum possible value $W_{\mathrm{max}}=0.3300$. This example illustrates that the precision of calibration of $\eta$ should be of the order of $0.01$.
As a final remark, we note that the spatial multiplexing utilized in Fig. 1 can be replaced by temporal multiplexing. This alternative approach requires only a single detector that is preceded by tunable losses [81, 82]. The probabilities $R_{1}$ and $R_{2}$ can be estimated from the measured probabilities of non-clicks of the detector for three levels of transmittances of the lossy channel: $\eta$, $\eta T$, and $\eta(1-T)$. For balanced setup with $T=\frac{1}{2}$ this reduces to just two different loss factors $\eta$ and $\eta/2$.
VII Conclusions